OpenStudy (anonymous):

rationalize the denominator of 4/5+the sq root of 2

OpenStudy (anonymous):

\[\frac{4}{5+\sqrt2}\]?

OpenStudy (anonymous):

rationalize by multiplying the numerator and denominator by the conjugate of the denominator the conjugate of \(a+\sqrt b\) is \(a-\sqrt b\) and this works because \[(a+\sqrt b)(a-\sqrt b)=a^2-b\] a number with no radical

OpenStudy (anonymous):

in your example it is \[\frac{4}{5+\sqrt2}\times \frac{5-\sqrt 2}{5-\sqrt2}\] \[=\frac{4(5-\sqrt 2)}{5^2-2}\] and you can clean it up a little after that

OpenStudy (anonymous):

thank you! i get it now

OpenStudy (anonymous):

also i keep getting the wrong answer on this: (7+sq root of 2)(8- 3 sq root 2)